OpenStudy (anonymous):

I am having trouble with this pointless Simpson's Rule...-_- Integral of (9-x^2) from 3 to 5. Please help me out!

OpenStudy (anonymous):

\[\int\limits_{3}^{5} 9-x ^{2}=[9x-\frac{x ^3}{3} ]_{3}^{5} \]\[= 9(3)-\frac{3^3}{3}-9(5)+\frac{3^5}{3} =27-9-45+81\]\[=54\]

OpenStudy (anonymous):

Thank you, however, I was having trouble understanding how to solve it using Simpson's Rule.

OpenStudy (anonymous):

All of my values were coming out negative. =(

OpenStudy (anonymous):

That's because the answer is -44/3

OpenStudy (anonymous):

The above formula needs the 5 and the 3 switched.

OpenStudy (anonymous):

Why is that BigFudge?

OpenStudy (anonymous):

That's how they're solved. The upper bound is plugged in first and the lower bound is plugged in second and subtracted from the first. http://tutorial.math.lamar.edu/Classes/CalcI/ComputingDefiniteIntegrals.aspx

OpenStudy (anonymous):

In wikipedia it says that the simpson rule is an approximation of a definite integral and the rule is \[\int_a^b f(x) dx \approx \frac{b-a}{6}\left[f(a)+4f\left(\frac{a+b}{2}\right)+f(b)\right]\] your \(f(x) = 9-x^2\), b = 5, a =3 and then you just put everything into the formula \[\int_3^5 9-x^2 dx \approx \frac{5-3}{6}\left[(9-3^2)+4\left(9-\left(\frac{3+5}{2}\right)^2\right)+(9-5^2)\right]\] \[ \int_3^5 9-x^2 dx \approx \frac{2}{6}\left[-28-16\right] \] \[ \int_3^5 9-x^2 dx \approx \frac{2}{6}\left[-44\right] \] \[ \int_3^5 9-x^2 dx \approx -\frac{44}{3} \]

OpenStudy (anonymous):

That is how they are solved in terms of just integration. Simpson's Rule is numerical integration and the rule is: Sn = \[\Delta x/3 (y0 + 4y1 + 2y2 +4y3 +... +2y(N-2) + 4y(N-1) + yN)\] where \[\Delta x = (b-a)/N \]

OpenStudy (anonymous):

Thank you plitter for straightening that out for me. It just seems weird that regular integration gives a positive answer but this one gives a negative one.

OpenStudy (anonymous):

Regular integration still gives you the same answer just through a different process.

OpenStudy (anonymous):

Dang, you're right. I guess I just integrated wrong in the first place. Thank you!